6.2.2 X is the Gaussian (0,1) random var- iable. Find the CDF of Y = |X| and its expected value E[Y].
Answer #1

Solution x is a Gaussian (0,1) Codil y=1x1. Fyly/= Plysy] = PC Y <IX!) = P(-y SX cy) P (x< 3) - P(X < - 3) = Exly) - Ex(Y) FyDifferentiate Cidif win to y gives as b dif. tylol = txly) + fxl-y) f(3) = fxly) ~ N101) is txly) - fxl-Y) even function) typNow E = Jo Jt A lette at it - La te at x-2 [e-c] = 12 (1) LEH = 1thank you